# Closed-Form Feedback-Free Learning with Forward Projection

### v3

@@ A.1 Local Learning Methods
Figure 5: Training Procedures for Forward Projection, Local Supervision, Forward Forward and Backpropagation learning algorithms for the $l$-th hidden layer. A: Forward Projection generates target matrix $\hat{\textbf{Z}}_{l}$ by projecting pre-synaptic inputs $\textbf{A}_{l-1}$ by $\textbf{Q}_{l}$ and labels Y by $\textbf{U}_{l}$. Weights $\textbf{W}_{l}$ are fitted by regression, generating membrane potential $\textbf{Z}_{l}=\textbf{A}_{l-1}\textbf{W}_{l}$. B: In Local Supervision, an auxiliary prediction $\hat{\textbf{Y}}_{l}$ is generated as a projection of the post-synaptic outputs $\textbf{A}_{l}$, and $\textbf{W}_{l}$ is updated by a short backward pass (red arrows). C: In Forward-Forward, “positive” and “negative” pre-synaptic activities, ${\textbf{A}}_{l}$ and $\check{\textbf{A}}_{l}$, are generated from true and spurious data-label pairs, respectively. $\textbf{W}_{l}$ is updated to maximise positive activity whilst minimising negative activity. D: In backpropagation, $\textbf{W}_{l}$ is updated along its gradient with respect to the backpropagated error.
@@ A.2 Convolutional Layer Implementation for FP
We consider implementation of Forward Projection in a 2D convolutional layer over pre-synaptic input $\textbf{a}_{l-1}\in\mathbb{R}^{1\times m_{l-1}\times R\times C}$, having $m_{l-1}$ channels, and spatial dimensions $R\times C$ representing rows and columns. The $k_{1}\times k_{2}$ convolutional kernel positioned in row $r\in\{1,\ldots,R\}$ and column $c\in\{1,\ldots,C\}$ is represented in “flattened” form as $\textbf{a}_{l-1,r,c}\in\mathbb{R}^{1\times(m_{l-1}k_{1}k_{2})}$. This pre-synaptic input is projected by the constant matrix $\textbf{Q}_{l}\in\mathbb{R}^{(m_{l-1}k_{1}k_{2})\times m_{l}}$ to generate local target potentials $\tilde{\textbf{z}}_{l,r,c}\in\mathbb{R}^{1\times m_{l}}$, such that
$\tilde{\textbf{z}}_{l,r,c}=g_{l}\left(\textbf{a}_{l-1,r,c}\textbf{Q}_{l}\right)+g_{l}\left(\textbf{yU}_{l}\right).$ (6)
@@
$\textbf{W}_{l}:=\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\textbf{A}_{l-1,r,c}^{T}\textbf{A}_{l-1,r,c}+\lambda\textbf{I}\right)^{-1}\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\textbf{A}_{l-1,r,c}^{T}\tilde{\textbf{Z}}_{l,r,c}\right).$ (7)
Here, $\textbf{A}_{l-1,r,c}\in\mathbb{R}^{N\times(m_{l-1}k_{1}k_{2})}$ is a matrix representing the pre-synaptic input for the convolutional kernel in row $r$ and column $c$ of the $l$-th layer over all $N$ training instances. Likewise, $\tilde{\textbf{Z}}_{l,r,c}\in\mathbb{R}^{N\times m_{l}}$ contains target potentials for the kernel in row $r$ and column $c$ of the $l$-th layer over all $N$ training instances. In each task, the FP penalty parameter was fixed at $\lambda=10$ for hidden layers and $\lambda=1$ for output layers. To prevent numerical overflow due to large layers or training datasets, a scalar reduction factor $\tau\in(0,1]$may be applied to downscale $\textbf{A}^{T}\textbf{A}$, $\lambda$, and $\textbf{A}^{T}\textbf{Z}$ as necessary, such that:
$\textbf{W}_{l}:=\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\tau\textbf{A}_{l-1,r,c}^{T}\textbf{A}_{l-1,r,c}+\tau\lambda\textbf{I}\right)^{-1}\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\tau\textbf{A}_{l-1,r,c}^{T}\tilde{\textbf{Z}}_{l,r,c}\right).$ (8)
@@ A.3 Attention Layer Implementation for FP
We consider the implementation of Forward Projection in a transformer layer over pre-synaptic input $\textbf{a}_{l-1}\in\mathbb{R}^{1\times R\times m_{l-1}}$, with $R$ patches, each having $m_{l-1}$ channels. Each attention head of dimension $d\in\mathbb{N}$ contains three weight matrices for optimisation: $\textbf{W}_{l}^{\operatorname{Query}}$, $\textbf{W}_{l}^{\operatorname{Key}}$ and $\textbf{W}_{l}^{\operatorname{Value}}$, yielding intermediate activations $\textbf{z}_{l}^{\operatorname{Query}}=\textbf{a}_{l-1}\textbf{W}_{l}^{\operatorname{Query}}$, $\textbf{z}_{l}^{\operatorname{Key}}=\textbf{a}_{l-1}\textbf{W}_{l}^{\operatorname{Key}}$, and $\textbf{z}_{l}^{\operatorname{Value}}=\textbf{a}_{l-1}\textbf{W}_{l}^{\operatorname{Value}}$ respectively. The attention scores $\textbf{s}_{l}$ are then calculated as
$\textbf{s}_{l}:={\operatorname{Softmax}\left({\frac{\textbf{z}_{l}^{\operatorname{Query}}\times(\textbf{z}_{l}^{\operatorname{Key}})^{T}}{\sqrt{d}}}\right)\times\textbf{z}_{l}^{\operatorname{Value}}}$ (9)
Forward projection may be applied to generate target values for each of $\textbf{z}_{l}^{\operatorname{Query}},\textbf{z}_{l}^{\operatorname{Key}},\textbf{z}_{l}^{\operatorname{Value}}$, using distinct data projection matrices $\textbf{Q}_{l}^{\operatorname{Query}},\textbf{Q}_{l}^{\operatorname{Key}},\textbf{Q}_{l}^{\operatorname{Value}}$, and a shared label projection matrix $\textbf{U}_{l}$, such that for each patch, query matrix target potentials are given by
$\tilde{\textbf{z}}_{l,r}^{\operatorname{Query}}=g_{l}\left(\textbf{a}_{l-1,r}\textbf{Q}_{l}^{\operatorname{Query}}\right)+g_{l}\left(\textbf{yU}_{l}\right).$ (10)
Thus, the query weight matrix is given by the penalised regression solution. Weights were fitted such that
$\textbf{W}_{l}^{\operatorname{Query}}:=\left(\sum^{R}_{r=1}\textbf{A}_{l-1,r}^{T}\textbf{A}_{l-1,r}+\lambda\textbf{I}\right)^{-1}\left(\sum^{R}_{r=1}\textbf{A}_{l-1,r}^{T}\tilde{\textbf{Z}}_{l,r}^{\operatorname{Query}}\right).$ (11)
$\textbf{W}_{l}^{\operatorname{Key}}$ and $\textbf{W}_{l}^{\operatorname{Value}}$, are fitted in the same manner - using $\textbf{Q}_{l}^{\operatorname{Key}}$ and $\textbf{Q}_{l}^{\operatorname{Value}}$, respectively, in place of $\textbf{Q}_{l}^{\operatorname{Query}}$ in (10).
@@ A.4 Implementation of local learning methods
Local Supervision was implemented by generating auxiliary predictions $\hat{\textbf{y}}_{1},\ldots,\hat{\textbf{y}}_{L-1}$ from a set of fixed linear operators $\textbf{U}_{1},\ldots\textbf{U}_{L-1}$ such that $\hat{\textbf{y}}_{l}=f_{L}(\textbf{a}_{l}\textbf{U}_{l})$. In each layer, $\textbf{W}_{l}$ was optimised to minimise the auxiliary loss $\mathcal{L_{l}}(\hat{\textbf{y}}_{l},\textbf{y})$ via gradient descent with two layers of backpropagation, such that
$\nabla_{\textbf{W}_{l}}\mathcal{L}_{l}=f_{l}^{\prime}(\textbf{U}_{l}^{T})\cdot f_{L}^{\prime}(\nabla_{\hat{\textbf{y}}_{l}}\mathcal{L}_{l})\cdot\textbf{a}^{T}_{l-1}.$ (12)
@@
$\hat{\textbf{y}}_{l}=\frac{1}{RC}\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\textbf{a}_{l-1,r,c}^{T}\textbf{a}_{l-1,r,c}\right)\textbf{U}_{l}.$ (13)
The Forward-Forward algorithm was implemented by generating both “positive” data-label pairs $[x,y]$ and “negative” data pairs $x,\check{y}]$ by concatenation. Positive activations were generated such that
$\textbf{a}_{l}=f_{l}(\textbf{W}_{l}\ldots f_{1}(\textbf{W}_{1}[\textbf{x},\textbf{y}])\ldots).$ (14)
@@
$\check{\textbf{a}}_{l}=f_{l}(\textbf{W}_{l}\ldots f_{1}(\textbf{W}_{1}[\textbf{x},\check{\textbf{y}}])\ldots).$ (15)
Thus, local auxiliary loss functions were computed using threshold hyperparameter $\theta=2$ and logistic sigmoid function $\sigma$, computing auxiliary loss $\mathcal{L}_{l}$ the such that
$\mathcal{L}_{l}:=\sigma(\lVert\tilde{\textbf{a}}_{l}\rVert^{2}_{2}-\theta)+\sigma(\theta-\lVert\textbf{a}_{l}\rVert^{2}_{2}).$ (16)
In MLP implementations, concatenation was implemented with supplementary input neurons to hold the label information as a one-hot vector. In convolutional implementations, the label was concatenated in the channel dimension, such that each channel indicated a single class, with a constant value over all kernel positions. Predictive Coding [8, 5] was implemented using inverse layers $\textbf{W}_{l}^{\operatorname{Back}}\in\mathbb{R}^{m_{l}\times m_{l-1}}$ for each weight matrix $\textbf{W}_{l}$, to estimate the pre-synaptic inputs from the post-synaptic outputs, such that
$\hat{\textbf{a}}_{l-1}:=f_{l}^{\operatorname{Back}}\left(\textbf{a}_{l}\textbf{W}_{l}^{\operatorname{Back}}\right)$ (17)
@@
$\mathcal{L}_{l}:=\lVert\textbf{a}_{l-1}-\hat{\textbf{a}}_{l-1}\rVert_{2}^{2}$ (18)
@@
$\hat{\textbf{a}}_{l-1}:=\textbf{a}_{l-1}-f_{l}^{\operatorname{Back}}\left(\textbf{a}_{l}\textbf{W}_{l}^{\operatorname{Back}}\right)+f_{l}^{\operatorname{Back}}\left(\hat{\textbf{a}}_{l}\textbf{W}_{l}^{\operatorname{Back}}\right)$ (19)
In our implementations, we set $f_{l}^{\operatorname{Back}}=\operatorname{ReLU}$ in both Predictive Coding and Difference Target Propagation models.
@@ A.5 Analysis of label projection approaches
We now consider the necessity of pre-synaptic input projections, which distinguish Forward Projection from alternative feedback-free approaches such as simple label projection and noisy label projection. The objective is to generate a weight matrix $\textbf{W}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$, given pre-synaptic inputs, $\textbf{A}_{l-1}\in\mathbb{R}^{N\times m_{l-1}}$, and label matrix $\textbf{Y}\in\mathbb{R}^{N\times m_{L}}$. As described previously, given some target potentials $\tilde{\textbf{Z}}$, weights will be fitted by equation (2), such that
$\textbf{W}_{l}=(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}+\lambda\textbf{I})^{-1}\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}$. (20)
@@
We first consider target generation by simple label projection, such that $\tilde{\textbf{Z}}_{l}:=\textbf{YU}_{l}$. Here we have $\operatorname{rank}(\tilde{\textbf{Z}})\leq\operatorname{rank}(\textbf{Y})\leq m_{L}$. Where $m_{l}>>m_{L}$, this implies severe degeneracy of $\textbf{W}_{l}$, predisposing to correlated or redundant neuronal activities in $\textbf{A}_{l}$. We next consider target generation by noisy label projection, using random Gaussian noise matrix $\textbf{E}\in\mathbb{R}^{N\times m_{l}}$ to perturb targets such that $\tilde{\textbf{Z}}_{l}:=\textbf{Y}\textbf{U}_{l}+\textbf{E}$. In this case, we have $\operatorname{rank}(\tilde{\textbf{Z}}_{l})\leq\operatorname{rank}(\textbf{E})\leq m_{l}$. However, as E is random and independent of Y, this perturbation strategy may adversely affect label modelling of Y. By adding projections of label and pre-synaptic inputs, such that $\tilde{\textbf{Z}}_{l}:=\textbf{a}_{l-1}\textbf{Q}_{l}+\textbf{Y}\textbf{U}_{l}$, we improve the upper bound to
$\operatorname{rank}(\tilde{\textbf{Z}}_{l})\leq\operatorname{rank}(\textbf{A}_{l-1})+\operatorname{rank}(\textbf{Y})\leq m_{l-1}+m_{L},$ (21)
without fitting $\textbf{W}_{l}$ to noise. Accordingly, the targets generated by our proposed method $\tilde{\textbf{Z}}_{l}:=g_{l}(\textbf{a}_{l-1}\textbf{Q}_{l})+g_{l}(\textbf{Y}\textbf{U}_{l})$ are subject to the same upper bound $\operatorname{rank}(\tilde{\textbf{Z}}_{l})\leq(m_{l-1}+m_{L})$.

### v4

@@ A.1 Local Learning Methods
Figure 5: Training Procedures for Forward Projection, Local Supervision, Forward Forward and Backpropagation learning algorithms for the $l$-th hidden layer. A: Forward Projection generates target matrix $\hat{\mathbf{Z}}_{l}$ by projecting pre-synaptic inputs $\mathbf{A}_{l-1}$ by $\mathbf{Q}_{l}$ and labels $\mathbf{Y}$ by $\mathbf{U}_{l}$. Weights $\mathbf{W}_{l}$ are fitted by regression, generating membrane potential $\mathbf{Z}_{l}=\mathbf{A}_{l-1}\mathbf{W}_{l}$. B: In Local Supervision, an auxiliary prediction $\hat{\mathbf{Y}}_{l}$ is generated as a projection of the post-synaptic outputs $\mathbf{A}_{l}$, and $\mathbf{W}_{l}$ is updated by a short backward pass (red arrows). C: In Forward-Forward, “positive” and “negative” pre-synaptic activities, ${\mathbf{A}}_{l}$ and $\check{\mathbf{A}}_{l}$, are generated from true and spurious data-label pairs, respectively. $\mathbf{W}_{l}$ is updated to maximise positive activity whilst minimising negative activity. D: In backpropagation, $\mathbf{W}_{l}$ is updated along its gradient with respect to the backpropagated error.
@@ A.2 Convolutional Layer Implementation for FP
We consider implementation of Forward Projection in a 2D convolutional layer over pre-synaptic input $\mathbf{a}_{l-1}\in\mathbb{R}^{1\times m_{l-1}\times R\times C}$, having $m_{l-1}$ channels, and spatial dimensions $R\times C$ representing rows and columns. The $k_{1}\times k_{2}$ convolutional kernel positioned in row $r\in\{1,\ldots,R\}$ and column $c\in\{1,\ldots,C\}$ is represented in “flattened” form as $\mathbf{a}_{l-1,r,c}\in\mathbb{R}^{1\times(m_{l-1}k_{1}k_{2})}$. This pre-synaptic input is projected by the constant matrix $\mathbf{Q}_{l}\in\mathbb{R}^{(m_{l-1}k_{1}k_{2})\times m_{l}}$ to generate local target potentials $\tilde{\mathbf{z}}_{l,r,c}\in\mathbb{R}^{1\times m_{l}}$, such that
$\tilde{\mathbf{z}}_{l,r,c}=g_{l}\left(\mathbf{a}_{l-1,r,c}\mathbf{Q}_{l}\right)+g_{l}\left(\mathbf{yU}_{l}\right).$ (6)
@@
$\mathbf{W}_{l}:=\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\mathbf{A}_{l-1,r,c}^{\top}\mathbf{A}_{l-1,r,c}+\lambda\mathbf{I}\right)^{-1}\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\mathbf{A}_{l-1,r,c}^{\top}\tilde{\mathbf{Z}}_{l,r,c}\right).$ (7)
Here, $\mathbf{A}_{l-1,r,c}\in\mathbb{R}^{N\times(m_{l-1}k_{1}k_{2})}$ is a matrix representing the pre-synaptic input for the convolutional kernel in row $r$ and column $c$ of the $l$-th layer over all $N$ training instances. Likewise, $\tilde{\mathbf{Z}}_{l,r,c}\in\mathbb{R}^{N\times m_{l}}$ contains target potentials for the kernel in row $r$ and column $c$ of the $l$-th layer over all $N$ training instances. In each task, the FP penalty parameter was fixed at $\lambda=10$ for hidden layers and $\lambda=1$ for output layers. To prevent numerical overflow due to large layers or training datasets, a scalar reduction factor $\tau\in(0,1]$ may be applied to downscale $\mathbf{A}^{\top}\mathbf{A}$, $\lambda$, and $\mathbf{A}^{\top}\mathbf{Z}$ as necessary, such that:
$\mathbf{W}_{l}:=\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\tau\mathbf{A}_{l-1,r,c}^{\top}\mathbf{A}_{l-1,r,c}+\tau\lambda\mathbf{I}\right)^{-1}\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\tau\mathbf{A}_{l-1,r,c}^{\top}\tilde{\mathbf{Z}}_{l,r,c}\right).$ (8)
@@ A.3 Attention Layer Implementation for FP
We consider the implementation of Forward Projection in a transformer layer [44] over pre-synaptic input $\mathbf{a}_{l-1}\in\mathbb{R}^{R\times m_{l-1}}$, with $R$ patches, each having $m_{l-1}$ channels. Each attention head of dimension $d\in\mathbb{N}$ contains three weight matrices for optimisation: $\mathbf{W}_{l}^{\operatorname{Query}}$, $\mathbf{W}_{l}^{\operatorname{Key}}$ and $\mathbf{W}_{l}^{\operatorname{Value}}$, yielding intermediate activations $\mathbf{z}_{l}^{\operatorname{Query}}=\mathbf{a}_{l-1}\mathbf{W}_{l}^{\operatorname{Query}}$, $\mathbf{z}_{l}^{\operatorname{Key}}=\mathbf{a}_{l-1}\mathbf{W}_{l}^{\operatorname{Key}}$, and $\mathbf{z}_{l}^{\operatorname{Value}}=\mathbf{a}_{l-1}\mathbf{W}_{l}^{\operatorname{Value}}$ respectively. The attention outputs $\mathbf{s}_{l}$ are then calculated as
$\mathbf{s}_{l}:={\operatorname{Softmax}\left({\frac{\mathbf{z}_{l}^{\operatorname{Query}}\times\left(\mathbf{z}_{l}^{\operatorname{Key}}\right)^{\top}}{\sqrt{d}}}\right)\times\mathbf{z}_{l}^{\operatorname{Value}}}$ (9)
Forward projection may be applied to generate target values for each of $\mathbf{z}_{l}^{\operatorname{Query}},\mathbf{z}_{l}^{\operatorname{Key}},\mathbf{z}_{l}^{\operatorname{Value}}$, using distinct data projection matrices $\mathbf{Q}_{l}^{\operatorname{Query}},\mathbf{Q}_{l}^{\operatorname{Key}},\mathbf{Q}_{l}^{\operatorname{Value}}$, and a shared label projection matrix $\mathbf{U}_{l}$, such that for each patch, query matrix target potentials are given by
$\tilde{\mathbf{z}}_{l,r}^{\operatorname{Query}}=g_{l}\left(\mathbf{a}_{l-1,r}\mathbf{Q}_{l}^{\operatorname{Query}}\right)+g_{l}\left(\mathbf{yU}_{l}\right).$ (10)
Thus, the query weight matrix is given by the penalised regression solution [45]. Weights were fitted such that
$\mathbf{W}_{l}^{\operatorname{Query}}:=\left(\sum^{R}_{r=1}\mathbf{A}_{l-1,r}^{\top}\mathbf{A}_{l-1,r}+\lambda\mathbf{I}\right)^{-1}\left(\sum^{R}_{r=1}\mathbf{A}_{l-1,r}^{\top}\tilde{\mathbf{Z}}_{l,r}^{\operatorname{Query}}\right).$ (11)
$\mathbf{W}_{l}^{\operatorname{Key}}$ and $\mathbf{W}_{l}^{\operatorname{Value}}$, are fitted in the same manner - using $\mathbf{Q}_{l}^{\operatorname{Key}}$ and $\mathbf{Q}_{l}^{\operatorname{Value}}$, respectively, in place of $\mathbf{Q}_{l}^{\operatorname{Query}}$ in (10).
@@ A.4 Implementation of local learning methods
Local Supervision was implemented by generating auxiliary predictions $\hat{\mathbf{y}}_{1},\ldots,\hat{\mathbf{y}}_{L-1}$ from a set of fixed linear operators $\mathbf{U}_{1},\ldots\mathbf{U}_{L-1}$ such that $\hat{\mathbf{y}}_{l}=f_{L}(\mathbf{a}_{l}\mathbf{U}_{l}^{+})$. In each layer, $\mathbf{W}_{l}$ was optimised to minimise the auxiliary loss $\mathcal{L_{l}}(\hat{\mathbf{y}}_{l},\mathbf{y})$ via gradient descent with two layers of backpropagation, such that
$\nabla_{\mathbf{W}_{l}}\mathcal{L}_{l}=\mathbf{a}^{\top}_{l-1}\left(f_{l}^{\prime}(\mathbf{z}_{l})\odot\left(f_{L}^{\prime}(\mathbf{a}_{l}\mathbf{U}_{l}^{+})\odot\nabla_{\hat{\mathbf{y}}_{l}}\mathcal{L}_{l}\right)(\mathbf{U}_{l}^{+})^{\top}\right).$ (12)
@@
$\hat{\mathbf{y}}_{l}=f_{L}\left(\frac{1}{RC}\sum^{R}_{r=1}\sum^{C}_{c=1}\mathbf{a}_{l,r,c}\mathbf{U}_{l}^{+}\right).$ (13)
The Forward-Forward algorithm [16] was implemented by generating both “positive” data-label pairs $\mathbf{x},\mathbf{y}]$ and “negative” data pairs $\mathbf{x},\check{\mathbf{y}}]$ by concatenation. Positive activations were generated such that
$\mathbf{a}^{\text{pos}}_{l}=f_{l}(\mathbf{W}_{l}\ldots f_{1}(\mathbf{W}_{1}[\mathbf{x},\mathbf{y}])\ldots).$ (14)
@@
$\mathbf{a}^{\text{neg}}_{l}=f_{l}(\mathbf{W}_{l}\ldots f_{1}(\mathbf{W}_{1}[\mathbf{x},\check{\mathbf{y}}])\ldots).$ (15)
Thus, local auxiliary loss functions were computed using threshold hyperparameter $\theta=2$ and logistic sigmoid function $\sigma$, computing auxiliary loss $\mathcal{L}_{l}$ such that
$\mathcal{L}_{l}:=\sigma(\lVert\mathbf{a}^{\text{neg}}_{l}\rVert^{2}_{2}-\theta)+\sigma(\theta-\lVert\mathbf{a}^{\text{pos}}_{l}\rVert^{2}_{2}).$ (16)
In MLP implementations, concatenation was implemented with supplementary input neurons to hold the label information as a one-hot vector. In convolutional implementations, the label was concatenated in the channel dimension, such that each channel indicated a single class, with a constant value over all kernel positions. Predictive Coding [8, 5] was implemented using inverse layers $\mathbf{W}_{l}^{\text{back}}\in\mathbb{R}^{m_{l}\times m_{l-1}}$ for each weight matrix $\mathbf{W}_{l}$, to estimate the pre-synaptic inputs from the post-synaptic outputs, such that
$\hat{\mathbf{a}}_{l-1}:=f_{l}^{\text{back}}\left(\mathbf{a}_{l}\mathbf{W}_{l}^{\text{back}}\right)$ (17)
@@
$\mathcal{L}_{l}:=\lVert\mathbf{a}_{l-1}-\hat{\mathbf{a}}_{l-1}\rVert_{2}^{2}$ (18)
@@
$\hat{\mathbf{a}}_{l-1}:=\mathbf{a}_{l-1}-f_{l}^{\text{back}}\left(\mathbf{a}_{l}\mathbf{W}_{l}^{\text{back}}\right)+f_{l}^{\text{back}}\left(\hat{\mathbf{a}}_{l}\mathbf{W}_{l}^{\text{back}}\right)$ (19)
In our implementations, we set $f_{l}^{\text{back}}=\operatorname{ReLU}$ in both Predictive Coding and Difference Target Propagation models.
@@ A.5 Analysis of label projection approaches
We now consider the necessity of pre-synaptic input projections, which distinguish Forward Projection from alternative feedback-free approaches such as simple label projection and noisy label projection. The objective is to generate a weight matrix $\mathbf{W}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$, given pre-synaptic inputs, $\mathbf{A}_{l-1}\in\mathbb{R}^{N\times m_{l-1}}$, and label matrix $\mathbf{Y}\in\mathbb{R}^{N\times m_{L}}$. As described previously, given some target potentials $\tilde{\mathbf{Z}}$, weights will be fitted by equation (2), such that
$\mathbf{W}_{l}=(\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1}+\lambda\textbf{I})^{-1}\mathbf{A}_{l-1}^{\top}\tilde{\mathbf{Z}}$. (20)
@@
We first consider target generation by simple label projection, such that $\tilde{\mathbf{Z}}_{l}:=\mathbf{YU}_{l}$. Here we have $\operatorname{rank}(\tilde{\mathbf{Z}})\leq\operatorname{rank}(\mathbf{Y})\leq m_{L}$. Where $m_{l}\gg m_{L}$, this implies severe degeneracy of $\mathbf{W}_{l}$, predisposing to correlated or redundant neuronal activities in $\mathbf{A}_{l}$. We next consider target generation by noisy label projection, using random Gaussian noise matrix $\mathbf{E}\in\mathbb{R}^{N\times m_{l}}$ to perturb targets such that $\tilde{\mathbf{Z}}_{l}:=\mathbf{Y}\mathbf{U}_{l}+\mathbf{E}$. In this case, we have
$\operatorname{rank}(\tilde{\mathbf{Z}}_{l})\leq\operatorname{min}\left(m_{l},\left(\operatorname{rank}(\mathbf{Y})+\operatorname{rank}(\mathbf{E})\right)\right)=m_{l},$ (21)
However, as $\mathbf{E}$ is random and independent of $\mathbf{Y}$, this perturbation strategy may adversely affect label modelling of $\mathbf{Y}$. By adding projections of label and pre-synaptic inputs, such that $\tilde{\mathbf{Z}}_{l}:=\mathbf{A}_{l-1}\mathbf{Q}_{l}+\mathbf{Y}\mathbf{U}_{l}$, we improve the upper bound to
$\operatorname{rank}(\tilde{\mathbf{Z}}_{l})\leq\operatorname{min}\left(m_{l},\left(\operatorname{rank}(\mathbf{A}_{l-1})+\operatorname{rank}(\mathbf{Y})\right)\right)\leq\operatorname{min}\left(m_{l},(m_{l-1}+m_{L})\right),$ (22)
without fitting $\mathbf{W}_{l}$ to noise. Accordingly, the targets generated by our proposed method $\tilde{\mathbf{Z}}_{l}:=g_{l}(\mathbf{a}_{l-1}\mathbf{Q}_{l})+g_{l}(\mathbf{Y}\mathbf{U}_{l})$ may also have full rank.